# AnyCap Project: A Unified Framework, Dataset, and Benchmark for Controllable Omni-modal Captioning

### v1

@@ 5 Conclusion
We present AnyCapModel, a lightweight and plug-and-play framework enhancing caption controllability across image, video, and audio modalities, without the costly fine-tuning of base models. Motivated by the scarcity of controllable caption data, we introduce a large-scale omni-modal caption preference dataset spanning three modalities and 28 distinct control dimensions. Furthermore, to our best of knowledge, we establish AnyCapEval, the first benchmark designed to rigorously assess the controllability of captions. Together, these contributions offer a unified solution to controllable captioning, bridging model design, data, and evaluation.
Limitations and future work. Our approach is inherently extensible and could readily generalize to emerging modalities such as 3D structures or molecular data. However, the scarcity of high-quality caption datasets in these domains, coupled with prohibitive human annotation costs, limits immediate applicability. Future availability of suitable datasets could unlock wider applications of our method.
Ethics and impact. While AnyCapDataset is carefully curated to support fine-grained controllability, there remains a risk of malicious misuse (for instance, deliberately creating harmful or misleading Q-A pairs). Models trained on such compromised data could exhibit significantly degraded controllability, resulting in severe negative impacts. We urge caution and responsible practices when utilizing or expanding upon our resources.
@@ A.2 The Singularity and Scarcity of User Instructions
Most classic datasets, such as COCO [69] and Flickr30k [70] in the image domain, MSR-VTT [71] and MSVD [72] in the video domain, and AudioCaps [61] and Clotho [65] in the audio domain, primarily provide “media–caption pairs”. They typically include multiple reference captions provided by different annotators. This “multi-reference” design is mainly intended to evaluate the diversity and coverage of descriptions, rather than providing explicit control during generation. The model’s primary task is to generate a “reasonable” description, but the user cannot specify requirements at generation time (e.g., “generate a humorous caption”, “describe the background elements in detail”). The only implicit user instruction is essentially “describe this content in brief”.
@@
Some datasets offer a degree of control, but often only along a single dimension. Dense captioning datasets, like DCI [58] for images or video datasets providing temporal information [73], associate descriptions with specific image regions or video segments. The “region/time segment” here can be viewed as a user instruction, i.e., “describe this part”. However, this is usually the sole dimension of control. Datasets like ASD v2 [56] focus on describing relationships between entities in an image. The user instruction is the “entity pair” whose relationship needs to be described. Datasets guided by concepts like MiraData [74] or InstanceCap [60] require the output to follow a specific structure, which constitutes a form of control over the output format. Although such structured outputs often encompass several instruction types (e.g., Background, Detail), the formulation and granularity of these user instructions are typically fixed, lacking the flexibility for customized control requests tailored to each individual image, video or audio.
@@
Figure 5: ACM enables controllable captioning across modalities by refining base model outputs to better align with user instructions. Given a user instruction, it takes initial captions from a foundation model and corrects instruction violations (highlighted in red), producing compliant, instruction-following outputs (green) — all without requiring fine-tuning of the base model.
@@
In summary, the majority of existing public datasets lack the capability for flexible, multi-dimensional control over the generated content, thereby limiting the development and evaluation of controllable captioning technologies. Even when leveraging large language models such as GPT for annotation, the resulting captions often reflect general patterns and lack explicit controllability along desired dimensions.
@@
To address this gap, we construct AnyCapDataset(ACD), a large-scale dataset featuring diverse user instructions paired with high-quality compliant captions across multiple modalities. Leveraging this dataset, we train AnyCapModel(ACM) to achieve fine-grained controllable caption generation. As illustrated in Fig. 5, ACM can transform initial captions into targeted descriptions that precisely follow various user instructions, enabling nuanced and customizable outputs without requiring modifications to the base model architecture.
@@ A.3 Data Analysis
To construct ACD, we utilize multiple publicly available datasets spanning three modalities: images, videos, and audio. Specifically, we incorporate data from ASD v2, DCI, DOCCI, ShareGPT-4o, InstanceCap, and AudioCaps. From these datasets, we extract images, videos, and audio content, and employ various multimodal large language models(MLLMs) to generate diverse user instructions and captions based on predefined instructions.
@@
ACD encompasses three primary modalities: images, videos, and audio, comprising a total of 300k $(q,c,a)$ triplets. Among them, 125k belong to the image modality, 100k to the video modality, and 75k to the audio modality. Each modality includes multiple instruction types.
@@ 3. Focus on introducing appropriate levels of inaccuracy while maintaining plausibility
To generate the $(q,c,a)$ triplets in ACD, where $q$ denotes a user instruction, $c$ is a high-quality caption, and $a$ is a relatively suboptimal caption, we utilize several MLLMs guided by specifically designed instruction templates. The detailed templates employed in this work are presented in Table 12 and Table 13. These templates are constructed according to different instruction types to ensure structured generation and facilitate downstream analysis. Each raw data sample is paired with one or more tailored instructions to produce diverse outputs. The main MLLMs involved include GPT-4o mini, InternVL2.5-78B and InternVL2.5-8B, with GPT-4o contributing to a small subset. Specifically, most user instructions $q$ and high-quality captions $c$ are generated by InternVL2.5-78B, covering approximately 84% of the dataset; approximately 2% of image-based high-quality captions $c$ under instruction types such as Instance and Instance Appearance are generated by GPT-4o; about 5% of the video-based high-quality captions $c$ under the Movement instruction type originate from the InstanceCap dataset; and the audio modality data, comprising around 9%, are all generated using GPT-4o mini. The suboptimal captions $a$ are constructed in three ways: by InternVL2.5-8B based solely on $q$, by InternVL2.5-8B with hallucinations introduced through specific instructions, or by directly copying $c$, a strategy used by InternVL2.5-78B.
@@ B.1 Limitations of Existing Evaluation Methods
Metrics developed for machine translation—such as BLEU [16], ROUGE [75], METEOR [76], and CIDEr [17]—primarily focus on n-gram overlap or co-occurrence statistics. While they can measure fluency and lexical similarity, they fail to accurately capture semantic consistency. For instance, sentences like “I do love large models” and “I do not love large models” could receive similarly high scores, despite conveying opposite meanings. These metrics cannot effectively detect deep semantic divergence, factual errors, or noncompliance with user instructions.
@@
Some studies, including those utilizing benchmarks like [21] which aims for fine-grained descriptions, propose extracting key information points from captions using MLLMs and computing precision, recall, and F1 scores. However, keypoint extraction itself is unstable: the number and granularity of extracted points vary widely across runs. Our empirical analysis shows low correlation between model-extracted keypoints and human-annotated ones, suggesting that such extraction-based metrics poorly reflect the true information content of captions.
@@
Benchmarks such as VDC [77] highlight that QA-pair based evaluation suffers from instability, primarily due to the variable quality and granularity of the QA pairs themselves. Designing high-quality, representative QA items that faithfully capture fine-grained caption attributes remains an open challenge.
@@ B.2.1 Separation of Content and Style
This separation is crucial because content and style represent fundamentally different aspects of caption quality that require distinct evaluation approaches. Content adherence focuses on factual accuracy and instruction compliance: whether the caption correctly follows the user’s specified requirements about what to describe. This is objectively verifiable through user instruction alignment. In contrast, style evaluation assesses subjective qualities like fluency, coherence, and naturalness that reflect human-like expression. Merging these dimensions would conflate objective and subjective criteria, making it difficult to diagnose specific model weaknesses. The orthogonal evaluation allows for more precise identification of whether a model’s limitations lie in its ability to follow instructions (content) or in its linguistic quality (style), enabling targeted improvements to each aspect. Furthermore, different applications may prioritize these dimensions differently - some use cases demand strict content control while others emphasize stylistic quality, making separate evaluation essential for practical deployment decisions.
@@ B.2.2 Keypoint Density Metric
2. Semantic Similarity Allowed: Phrases with similar meaning should be considered matches (e.g., "holding a ball" and "grasping a sphere").
@@
{"caption_evaluation": {"key_points_scores": {"key_point_1": score, "key_point_2": score, …}, "total_score": sum_of_scores, "score_reasons": {"key_point_1": "reason for score", "key_point_2": "reason for score", …}}}
@@
{"caption_evaluation": {
"key_points_scores": {"mention the man’s position": 1, "describe the man’s appearance": 0, "mention the woman’s position": 1, "describe the woman’s appearance": 0, "mention the boy’s position": 1, "describe the boy’s action": 0},
"total_score": 3,
"score_reasons": {"mention the man’s position": "Correctly mentions standing on the left side of the table", "describe the man’s appearance": "Missing glasses reference", "mention the woman’s position": "Correctly mentions sitting on the right side of the table", "describe the woman’s appearance": "Missing red dress reference", "mention the boy’s position": "Correctly mentions crouching under the table", "describe the boy’s action": "Missing picking up a toy reference"}
@@
You are an expert evaluator tasked with scoring model outputs based on a specific rubric. Please carefully analyze the provided video frames, "Caption Type", "Model Output," and "Reference" text, and assign a score between 0 and 4 to "Model Output" using different criteria for different caption types. Ensure your scoring is consistent and strictly adheres to the definitions provided.
@@
- 1 (Poor): Significant quality issues OR major hallucination (>50% of the content is fictitious or contradictory).
- 2 (Below Average): Slightly inferior to reference OR limited hallucination (<50% of the content is inaccurate, but does not affect the core content).
@@
- High Quality: Length is within $\pm$30of the reference word count; concise and captures the core content of the video.
- Low Quality: Length exceeds $\pm$30of the reference word count; includes irrelevant details or omits key information.
@@
- High Quality: Length is within $\pm$30of the reference word count; provides rich descriptions of the video’s main elements, actions, and settings.
- Low Quality: Length exceeds $\pm$30of the reference word count; descriptions lack detail or include irrelevant information.
@@ 3. For caption quality, evaluate based on:
- For Caption Type that is brief or detail, ensure the model output’s word count is within $\pm$30of the reference word count. If not, the score cannot be higher than 1 for brief and detailed captions.
@@ 4. For hallucination, evaluate based on:
Mandatory Rule: For the Caption Type that is brief or detail, if the length exceeds $\pm$30of the reference word count, the score cannot be higher than 1.
@@ 6. Return your response in this format:
{"score": [0-4], "reason": "1-2 sentence explanation"}
@@ C.1 Experimental Setup
Table 16: Key hyperparameters used for training AnyCapModel.
@@
The experimental setup utilizes a computing infrastructure comprising 32 NVIDIA A100 GPUs, distributed across 4 nodes with 8 GPUs each. Each GPU is allocated 10 CPU cores, and all training jobs are managed using Slurm’s srun command with GPU reservation. For distributed training, we employ torch.distributed, launching the script via srun to ensure dynamic rank assignment and multi-node coordination. The model is trained using the AnyCapDataset. During training, the vision and audio backbones are kept frozen, while the MLP and LLM layers remain trainable. Image processing involves forcing the input image size to 448 pixels, with a maximum dynamic patch count of 6 and a down-sampling ratio of 0.5. Dynamic image sizing and thumbnail support are also enabled. The detailed hyperparameters used during training are listed in Table 16.
@@ C.2 Ablation Study
Effect of training objectives.
Regarding image modality, we compare AnyCapModel (ACM) with three widely used baseline methods: Supervised Fine-Tuning (SFT), Direct Preference Optimization (DPO), and Self-Critic (SC). Due to limitations in computational resources and budget, we conduct all comparative experiments using the lightweight InternVL2.5-2B on the MIA-Bench. As shown in the main paper, our method achieves a remarkable 16.9% improvement, significantly outperforming baseline methods.
C.3 Qualitative Examples for Main Experiments

### v2

@@ 5 Conclusion
We present AnyCap, a lightweight module that aligns base captioners with human instructions in omni-modal settings, enhancing instruction-following without retraining base models. It leverages residual correction in a unified multimodal design, enabling a single model to serve images, videos, and audio. To support this, we introduce AnyCapData, a large-scale, fine-grained instruction–based caption dataset. We also present AnyCapEval, a new benchmark to more accurately evaluate instruction alignment across modalities. On both public benchmarks and AnyCapEval, AnyCap consistently achieves superior performance, advancing instruction-aligned captioning across modalities.
Ethics statement. While AnyCapData is carefully curated to support fine-grained controllability, there remains a risk of malicious misuse (for instance, deliberately creating harmful or misleading Q-A pairs). Models trained on such compromised data could exhibit significantly degraded controllability, resulting in severe negative impacts. We urge caution and responsible practices when utilizing or expanding upon our resources.
Reproducibility statement. The framework and training strategy of AnyCap are described in Sec. 3.1, with detailed training settings and hyperparameters provided in Sec. C.1. The construction process, advantage comparison, data statistics, and quality assurance of AnyCapData are outlined in Sec. 3.2, with additional details and the required prompt templates provided in Appendix A. The testing process of AnyCapEval is fully described in Sec. 3.3. More details of AnyCapEval are discussed in Appendix B, including an analysis of its correlation with human judgments. The experimental setup and more detailed results for some experiments are provided in Appendix C, where we also include visualizations of the experimental outcomes and additional performance details of AnyCap on downstream tasks. All code and data will be open-sourced to ensure reproducibility and further research.
@@ A.2 The Singularity and Scarcity of User Instructions
Most classic datasets, such as COCO (Lin et al., 2014) and Flickr30k (Young et al., 2014) in the image domain, MSR-VTT (Xu et al., 2016) and MSVD (Chen & Dolan, 2011) in the video domain, and AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) in the audio domain, primarily provide “media–caption pairs”. They typically include multiple reference captions provided by different annotators. This “multi-reference” design is mainly intended to evaluate the diversity and coverage of descriptions, rather than providing explicit control during generation. The model’s primary task is to generate a “reasonable” description, but the user cannot specify requirements at generation time (e.g., “generate a humorous caption”, “describe the background elements in detail”). The only implicit user instruction is essentially “describe this content in brief”.
@@
Some datasets offer a degree of control, but often only along a single dimension. Dense captioning datasets, like DCI (Urbanek et al., 2024) for images or video datasets providing temporal information (Krishna et al., 2017a), associate descriptions with specific image regions or video segments. The “region/time segment” here can be viewed as a user instruction, i.e., “describe this part”. However, this is usually the sole dimension of control. Datasets like ASD v2 (Wang et al., 2024b) focus on describing relationships between entities in an image. The user instruction is the “entity pair” whose relationship needs to be described. Datasets guided by concepts like MiraData (Ju et al., 2024) or InstanceCap (Fan et al., 2024) require the output to follow a specific structure, which constitutes a form of control over the output format. Although such structured outputs often encompass several instruction types (e.g., Background, Detail), the formulation and granularity of these user instructions are typically fixed, lacking the flexibility for customized control requests tailored to each individual image, video or audio.
@@
Figure 6: AnyCap enables controllable captioning across modalities by refining base model outputs to better align with user instructions. Given a user instruction, it takes initial captions from a foundation model and corrects instruction violations (highlighted in red), producing compliant, instruction-following outputs (green), all without requiring fine-tuning of the base model.
@@
In summary, the majority of existing public datasets lack the ability to align captions with specific user instructions across multiple dimensions, limiting the development and evaluation of instruction-aligned captioning technologies. Even when leveraging large language models such as GPT for annotation, the resulting captions often reflect general patterns and fail to align precisely with the specified instructions along desired dimensions.
@@
To address this gap, we construct AnyCapData, a large-scale dataset featuring diverse user instructions paired with high-quality, instruction-compliant captions across multiple modalities. Leveraging this dataset, we train AnyCapModel to achieve fine-grained instruction-aligned caption generation. As illustrated in Fig. 6, AnyCap can transform initial captions into highly specific descriptions that accurately follow various user instructions, enabling precise and customizable outputs without requiring modifications to the base model architecture.
@@ A.3 Data Analysis
To construct AnyCapData, we utilize multiple publicly available datasets spanning three modalities: images, videos, and audio. Specifically, we incorporate data from ASD v2, DCI, DOCCI, ShareGPT-4o, InstanceCap, and AudioCaps. From these datasets, we extract images, videos, and audio content, and employ various multimodal large language models (MLLMs) to generate diverse user instructions and captions based on predefined instructions.
@@
AnyCapData encompasses three primary modalities: images, videos, and audio, comprising a total of 300k $(q,c,a)$ triplets. Among them, 125k belong to the image modality, 100k to video modality, and 75k to audio modality. Each modality includes multiple instruction types.
@@ 3. Focus on introducing appropriate levels of inaccuracy while maintaining plausibility
To generate the $(q,c,a)$ triplets in AnyCapData, where $q$ denotes a user instruction, $c$ is a high-quality caption, and $a$ is a relatively suboptimal caption, we utilize several MLLMs guided by specifically designed instruction templates. The detailed templates employed in this work are presented in Table 14 and Table 15. These templates are constructed according to different instruction types to ensure structured generation and facilitate downstream analysis. Each raw data sample is paired with one or more tailored instructions to produce diverse outputs.
@@ B.1 Limitations of Existing Evaluation Methods
Metrics developed for machine translation, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), and CIDEr (Vedantam et al., 2015), primarily focus on n-gram overlap or co-occurrence statistics. While they can measure fluency and lexical similarity, they fail to accurately capture semantic consistency. For instance, sentences like “I do love large models” and “I do not love large models” could receive similarly high scores, despite conveying opposite meanings. These metrics cannot effectively detect deep semantic divergence, factual errors, or noncompliance with user instructions.
@@
Some studies, including those utilizing benchmarks like (Wang et al., 2024a) which aims for fine-grained descriptions, propose extracting key information points from captions using MLLMs and computing precision, recall, and F1 scores. However, keypoint extraction itself is unstable: the number and granularity of extracted points vary widely across runs. Our empirical analysis shows low correlation between model-extracted keypoints and human-annotated ones, suggesting that such extraction-based metrics poorly reflect the true information content of captions.
@@
Benchmarks such as VDC (Chai et al., 2024) highlight that QA-pair based evaluation suffers from instability, primarily due to the variable quality and granularity of the QA pairs themselves. Designing high-quality, representative QA items that faithfully capture fine-grained caption attributes remains an open challenge.
@@ B.2.1 Separation of Content and Style
This separation is crucial because content and style represent fundamentally different aspects of caption quality that require distinct evaluation approaches. Content adherence focuses on factual accuracy and instruction compliance: whether the caption correctly follows the user’s specified requirements about what to describe. This is objectively verifiable through user instruction alignment. In contrast, style evaluation assesses subjective qualities like fluency, coherence, and naturalness that reflect human-like expression. Merging these dimensions would conflate objective and subjective criteria, making it difficult to diagnose specific model weaknesses. The orthogonal evaluation allows for more precise identification of whether a model’s limitations lie in its ability to follow instructions (content) or in its linguistic quality (style), enabling targeted improvements to each aspect. Furthermore, different applications may prioritize these dimensions differently: some use cases demand strict content control while others emphasize stylistic quality, making separate evaluation essential for practical deployment decisions.
@@ B.2.2 Keypoint Density Metric
Correlation Type Info Points vs. Actual Info Caption Length vs. Actual Info
Pearson 0.284 0.521
Spearman 0.256 0.461
Kendall’s Tau 0.231 0.344
Table 16: Correlations Comparison
@@
2. Semantic Similarity Allowed: Phrases with similar meaning should be considered matches (e.g., ”holding a ball” and ”grasping a sphere”).
@@
{”caption_evaluation”: {”key_points_scores”: {”key_point_1”: score, ”key_point_2”: score, …}, ”total_score”: sum_of_scores, ”score_reasons”: {”key_point_1”: ”reason for score”, ”key_point_2”: ”reason for score”, …}}}
@@
{”caption_evaluation”: {
”key_points_scores”: {”mention the man’s position”: 1, ”describe the man’s appearance”: 0, ”mention the woman’s position”: 1, ”describe the woman’s appearance”: 0, ”mention the boy’s position”: 1, ”describe the boy’s action”: 0},
”total_score”: 3,
”score_reasons”: {”mention the man’s position”: ”Correctly mentions standing on the left side of the table”, ”describe the man’s appearance”: ”Missing glasses reference”, ”mention the woman’s position”: ”Correctly mentions sitting on the right side of the table”, ”describe the woman’s appearance”: ”Missing red dress reference”, ”mention the boy’s position”: ”Correctly mentions crouching under the table”, ”describe the boy’s action”: ”Missing picking up a toy reference”}
@@
You are an expert evaluator tasked with scoring model outputs based on a specific rubric. Please carefully analyze the provided video frames, ”Caption Type”, ”Model Output,” and ”Reference” text, and assign a score between 0 and 4 to ”Model Output” using different criteria for different caption types. Ensure your scoring is consistent and strictly adheres to the definitions provided.
@@
- 1 (Poor): Significant quality issues OR major hallucination (¿50% of the content is fictitious or contradictory).
- 2 (Below Average): Slightly inferior to reference OR limited hallucination (¡50% of the content is inaccurate, but does not affect the core content).
@@
- High Quality: Length is within $\pm$30% of the reference word count; concise and captures the core content of the video.
- Low Quality: Length exceeds $\pm$30% of the reference word count; includes irrelevant details or omits key information.
@@
- High Quality: Length is within $\pm$30% of the reference word count; provides rich descriptions of the video’s main elements, actions, and settings.
- Low Quality: Length exceeds $\pm$30% of the reference word count; descriptions lack detail or include irrelevant information.
@@ 3. For caption quality, evaluate based on:
- For Caption Type that is brief or detail, ensure the model output’s word count is within $\pm$30% of the reference word count. If not, the score cannot be higher than 1 for brief and detailed captions.
@@ 4. For hallucination, evaluate based on:
Mandatory Rule: For the Caption Type that is brief or detail, if the length exceeds $\pm$30% of the reference word count, the score cannot be higher than 1.
@@ 6. Return your response in this format:
{”score”: [0-4], ”reason”: ”1-2 sentence explanation”}
@@ C.1 Experimental Setup
Table 19: Key hyperparameters used for training AnyCap.
@@
The experimental setup utilizes a computing infrastructure comprising 32 NVIDIA A100 GPUs, distributed across 4 nodes with 8 GPUs each. Each GPU is allocated 10 CPU cores, and all training jobs are managed using Slurm’s srun with GPU reservation. For distributed training, we employ torch.distributed, launching via srun to enable dynamic rank assignment and multi-node coordination. We train AnyCapModel on AnyCapData with the AdamW optimizer for 3 epochs, a learning rate of $1\times 10^{-6}$, a cosine learning-rate schedule with a 0.03 warmup ratio, weight decay of 0.01, a global batch size of 256, and bfloat16 mixed precision. As the modality backbones we use InternVL (vision) and EAT (audio). InternVL’s open training and fine-tuning pipeline facilitates hyperparameter tuning, and EAT attains state-of-the-art results on public benchmarks, making both solid foundations for our experiments. During AnyCap training, the external modality backbones are frozen, while the AnyCap internal components, including its language model and MLP layers, are updated. For images, we enforce an input size of 448 pixels with a maximum dynamic patch count of 6 and a down-sampling ratio of 0.5; dynamic image sizing and thumbnail support are enabled. The detailed hyperparameters used during training are listed in Table 19.
@@ C.2 Ablation Study
Table 25: Instruction-based video captioning on AnyCapEval.
Model Pos. App. Act. Ins. Persp. Cam. Bkg. Evt. Avg.
Proprietary Models
GPT-4o 2.41 4.00 3.86 4.01 3.74 2.70 4.66 3.03 3.55
+AnyCap-2B 3.45 (+1.04) 6.15 (+2.15) 6.26 (+2.40) 5.45 (+1.44) 4.58 (+0.84) 6.53 (+3.83) 5.39 (+0.73) 4.60 (+1.57) 5.30 (+1.75)
+AnyCap-8B 4.92 (+2.51) 6.60 (+2.60) 7.68 (+3.82) 4.95 (+0.94) 5.26 (+1.52) 5.67 (+2.97) 6.04 (+1.38) 4.81 (+1.78) 5.74 (+2.19)
Open-source Models
InternVL2.5-8B 3.08 4.33 3.44 3.78 3.26 3.16 4.56 2.55 3.52
+AnyCap-2B 4.67 (+1.59) 6.32 (+1.99) 3.19 (-0.25) 5.22 (+1.44) 4.66 (+1.40) 6.49 (+3.33) 5.49 (+0.93) 4.50 (+1.95) 5.07 (+1.55)
+AnyCap-8B 4.95 (+1.87) 6.39 (+2.06) 7.03 (+3.59) 5.49 (+1.71) 5.85 (+2.59) 5.66 (+2.50) 5.49 (+0.93) 4.95 (+2.40) 5.73 (+2.21)
Qwen2.5VL-7B 2.88 4.20 3.29 4.02 3.98 2.24 4.77 2.93 3.54
+AnyCap-2B 3.39 (+0.51) 5.97 (+1.77) 6.23 (+2.94) 4.91 (+0.89) 5.23 (+1.25) 6.30 (+4.06) 5.51 (+0.74) 4.48 (+1.55) 5.25 (+1.71)
+AnyCap-8B 4.03 (+1.15) 6.15 (+1.95) 6.58 (+3.29) 4.81 (+0.79) 6.24 (+2.26) 6.04 (+3.80) 5.74 (+0.97) 4.82 (+1.89) 5.55 (+2.01)
Model Brf. Det. Sty. Poe. Nar. Avg.
Proprietary Models
GPT-4o 1.47 1.52 2.77 2.52 2.48 2.15
+AnyCap-2B 1.97 (+0.50) 1.78 (+0.26) 2.52 (-0.25) 2.60 (+0.08) 2.65 (+0.17) 2.30 (+0.15)
+AnyCap-8B 1.95 (+0.48) 1.82 (+0.30) 2.58 (-0.19) 2.50 (-0.02) 2.77 (+0.29) 2.32 (+0.17)
Open-source Models
InternVL2.5-8B 1.39 1.77 2.23 1.91 2.34 1.93
+AnyCap-2B 1.84 (+0.45) 1.80 (+0.03) 2.45 (+0.22) 2.48 (+0.57) 2.39 (+0.05) 2.19 (+0.26)
+AnyCap-8B 2.00 (+0.61) 1.88 (+0.11) 2.50 (+0.27) 2.20 (+0.29) 2.61 (+0.27) 2.24 (+0.31)
Qwen2.5VL-7B 1.63 1.65 2.55 2.40 2.29 2.10
+AnyCap-2B 1.66 (+0.03) 1.82 (+0.17) 2.48 (-0.07) 2.30 (-0.10) 2.55 (+0.26) 2.16 (+0.06)
+AnyCap-8B 1.92 (+0.29) 1.98 (+0.33) 2.65 (+0.10) 2.45 (+0.05) 2.55 (+0.26) 2.31 (+0.21)
C.3 Further Results for the Main Experiments
Table 26: Performance of additional models on the image modality of AnyCapEval.
[…]
LLaVA-7B 1.50 2.99 3.85 2.63 2.74 1.27 0.98 1.21 0.71 1.67 1.19
 +AnyCap-8B 3.26 4.72 5.23 4.56 4.44 2.27 1.98 2.56 2.90 2.74 2.45
MiniCPM-o 1.96 3.80 4.90 3.07 3.43 2.11 2.06 2.15 1.74 2.54 2.12
 +AnyCap-8B 2.31 4.40 5.00 2.98 3.67 2.36 2.14 2.73 2.33 2.74 2.44
YiVL-34B 1.38 1.85 1.74 1.73 1.67 0.91 0.59 0.85 0.55 1.04 0.81
 +AnyCap-8B 3.54 4.69 4.93 4.66 4.45 1.98 2.04 2.59 2.65 2.59 2.30
InternVL-38B 1.69 3.79 4.92 3.01 3.35 2.15 2.02 2.17 2.06 2.64 2.21
 +AnyCap-8B 3.31 4.90 5.00 4.29 4.38 2.38 2.15 2.72 2.63 2.82 2.54
Table 27: Performance of additional models on the video modality of AnyCapEval. Top: Content metrics; Bottom: Style metrics.
Model Content
[…]
Tarsier2-7B 2.50 3.39 3.80 3.95 2.49 1.10 4.26 4.18 2.83
 +AnyCap-8B 4.85 7.79 7.96 5.94 5.77 5.48 6.10 5.57 6.18
LLaVA-Video 2.77 5.89 2.84 3.67 3.10 1.34 6.15 2.65 3.55
 +AnyCap-8B 4.37 7.08 6.20 5.68 5.79 5.35 6.62 4.35 5.68
MiniCPM-o 2.68 3.59 3.31 4.21 4.16 1.76 4.62 1.76 3.26
 +AnyCap-8B 4.65 5.70 6.05 5.20 5.48 5.06 6.01 5.06 5.40
Model Style
[…]
Tarsier2-7B 0.63 2.08 2.12 0.42 2.45 1.51
 +AnyCap-8B 1.87 1.90 2.65 2.38 2.55 2.26
LLaVA-Video 1.61 1.19 3.36 2.01 2.70 2.13
 +AnyCap-8B 1.97 2.11 2.74 0.99 1.32 2.23
MiniCPM-o 1.29 1.40 2.30 1.73 2.48 1.84
 +AnyCap-8B 1.97 2.11 2.74 0.99 1.32 2.23
In the main paper we reported the performance of representative models on AnyCapEval. Here we provide a broader set of results ( Tables 25, 26 and 27). Note that Tarsier2-7B, while strong at detailed captioning, attains comparatively lower scores on our benchmark because it tends to produce long, highly detailed descriptions that do not strictly follow the given instructions, indicating limited controllability. By contrast, AnyCap makes explicit use of rich instruction signals and achieves substantially better controllability and keypoint coverage. Concretely, with AnyCap we observe point counts of Content: 6.18 and Style: 2.26, compared with GPT-4o’s Content: 3.55 and Style: 2.15 under the same protocol.